Savannah and her soccer team of 18 players are invited to the womans national game in Toronto. There are 72 seats available. Every regular priced ticket is valued at $150.00 each. However, Savannah's mother was friends with the box office guy, and received a discount of 25% off.What was the total amount of money saved(in $)?
My question:
What are two let statements, and 2 equations that represent this word problem?

Answer by stanbon(75887)

Savannah and her soccer team of 18 players are invited to the woman's national game in Toronto.
------------------------
There are 72 seats available.
------
Every regular priced ticket is valued at $150.00 each.
------
However, Savannah's mother was friends with the box office guy, and received a discount of 25% off.
-------
What was the total amount of money saved(in $)?
---
Money saved = 0.25*150 = $37.50
My question:
What are two let statements, and 2 equations that represent this word problem?
---
price per discounted ticket = 150-0.25*150 = 0.75*150 = $112.50
---
Cost for the 72 tickets = 112.5*72 = $8100.00
==========
Cheers,
Stan H.
==========